OpenStudy (anonymous):

You have a deck of 52 cards. Find the odds in favor of choosing a King. 1/12 1/52 1/26 1/13

OpenStudy (anonymous):

4 in 52 so 1/13

OpenStudy (anonymous):

i agree 1/13

OpenStudy (whpalmer4):

that's the probability of choosing a king, but in fact the odds has a different meaning! the probability here is 4 in 52, or 4/52 (= 1/13) the odds are the probability of something happening / the probability of it not happening probability of not drawing a king is 48 in 52, or 48/52 the ODDS of choosing a king are thus (4/52) / (48/52) = 4/48 = 1/12 A is the correct answer, not D. @cc3 @jakebarnes @bobsue

OpenStudy (anonymous):

did not know that thx!